OpenStudy (anonymous):

CHECK MY SUPER EASY WORK FOR A QUICK MEDAL!!! What is the amplitude, period, and phase shift of f(x) = -3 cos(4x + π) + 6?

OpenStudy (anonymous):

I already know the amplitude is -3 and the phase shift is -pi/4, how do I find the period?

OpenStudy (anonymous):

@mathmate @ganeshie8

OpenStudy (kropot72):

If y = A cos(Bx + C) + D the amplitude is |A|, the absolute value of A the period T is: \[\large T=\frac{2 \pi}{B}\] the phase shift is: \[\large \frac{C}{B}\] You need to look again at your answers of the amplitude and the phase shift.

OpenStudy (kropot72):

answers for*

OpenStudy (anonymous):

@kropot72 okay so the answer for this would be amp = 3, per = pi / 2, and ps = -pi / 4, right?

OpenStudy (anonymous):

can someone check this^, please?

zepdrix (zepdrix):

ya looks good c:

OpenStudy (anonymous):

thanks!!!

OpenStudy (kropot72):

The phase shift is: \[\large \frac{\pi}{4}\]